OpenStudy (anonymous):

Directions: Solve equation by completing the square. r^2-4r-91=7

OpenStudy (zarkon):

\[r^2-4r-91=7\] add 91 \[r^2-4r=98\] take 1/2 of -4 and then square [\frac{1}{2}\times(-4)=-2\] \[(-2)^2=4\] add 4 to both sides \[r^2-4r+4=98+4\] \[r^2-4r+4=102\] now the left hand side factors \[(r-2)^2=102\] take square root of both sides \[r-2=\pm\sqrt{102}\] add 2 \[r=2\pm\sqrt{102}\]

OpenStudy (zarkon):

let me know if you have any other problems

OpenStudy (zarkon):

the part above that looks a little weird says \[\frac{1}{2}\times(-4)=-2\]